Hamid Oil (HOSB) is a sole proprietor, producing oil palm in a large scale. The land and machinery he
owns has a current market value of RM4 million. HOSB owes a local bank RM3 million. Last year
HOSB sold RM5 million worth of oil palm. Its variable operating cost were RM4.5 million; accounting
BMME5103/USTY/JAN16/A-KK
depreciation was RM40,000, although the actual decline in value of HOSB macinery was RM60,000
last year. Hamid paid himself a salary of RM50,000, which is not considered as part of HOSB variable
operating costs. Interest on the bank loan was RM400,000. If Hamid worked for another oil palm
producer or a local manufacturer, his annual income would be about RM30,000. Hamid can invest
any funds that would be derived, if the farm were sold, to earn 10 percent annually. (Ignore taxes)
a. Compute Hamid’s accounting profits.
b. Compute Hamid’s economics profits.
c. Based on the answers above, what would be the better option for Hamid, financially?
1
Expert's answer
2016-05-30T09:38:03-0400
The land and machinery - RM4 million,
loan in local bank - RM3 million,
goods sold - RM5 million,
variable operating cost - RM4.5 million;
accounting depreciation - RM40,000,
actual decline in value of HOSB macinery - RM60,000,
Hamid paid himself a salary of RM50,000 (part of variable operating costs).
Interest on the loan - RM400,000.
Hamid opportunity costs - RM30,000. Hamid can invest any funds that would be derived, if the farm were sold, to earn 10 percent annually.
a. Hamid’s accounting profits are:
Accounting profit = TR - TC = 5,000,000 - (4,500,000 +400,000 + 50,000 + (60,000 - 40,000)) = RM30,000.
b. Hamid’s economics profits are:
Economic profit = Accounting profit - Internal (opportunity costs) = 30,000 - 30,000 - 4,000,000*0.1 = -RM400,000.
c. Based on the answers above, the better option for Hamid financially is to continue producing until the loan is repaid.